OpenStudy (anonymous):

P.E.M.D.A.S. 3-3+9+90/10=3x8=34?? Correct?

OpenStudy (anonymous):

are you missing some brackets?

OpenStudy (anonymous):

The exact result would be 18, you must be missing some brackets

OpenStudy (anonymous):

3x8=24 btw

OpenStudy (anonymous):

No Brackets Missing Here's the problem Again I accidently typed it wrong sorry about that 3-3+9+90/10+3x8 My final answer should be 34

OpenStudy (anonymous):

answer is 42

OpenStudy (anonymous):

That should help

OpenStudy (anonymous):

3-3+9+90/10+3x8 =3-3+9+90/10+24 =3-3+9+9+24 =9+9+24 =18+24 =42

OpenStudy (anonymous):

Ok thank you guys

OpenStudy (anonymous):

No problem :)